OpenStudy (anonymous):

Which of the following is the equation of the horizontal asymptote of the graph of f(x)=(1/3) x+2^-3.

OpenStudy (anonymous):

\[f(x)=\left(\begin{matrix}1 \\ 3\end{matrix}\right)x+2\]

OpenStudy (anonymous):

-3 at the end.

OpenStudy (anonymous):

Yay! @apoorvk (:

OpenStudy (anonymous):

C'mon @campbell_st (:

OpenStudy (anonymous):

Like it has x+2 a little higher then -3

OpenStudy (anonymous):

@knock can help, too. (:

OpenStudy (anonymous):

It's minus 3. Not negative.

OpenStudy (campbell_st):

well to me its a linear equation if its \[y = \frac{1}{3} x + 2^{-3}\] as this is the same as \[y = \frac{1}{3} x + \frac{1}{8}\]

OpenStudy (anonymous):

No @campbell_st . Yes @apoorvk :D

OpenStudy (apoorvk):

@campbell_st she's new, doesn't know LaTeX yet..

OpenStudy (anonymous):

LaTeX?

OpenStudy (anonymous):

Nevermind lol

OpenStudy (apoorvk):

It's a syntax to write out equations the way we do. Let me see if I can find you a link to understand it, it's easy barely takes 2 minutes.

OpenStudy (campbell_st):

this an exponetial curve that hae been moved down 3 units... so the horizontal asymptote is y = -3

OpenStudy (anonymous):

I'm confused.. I just want what you were doing before @apoorvk

OpenStudy (apoorvk):

okay similar curve, but this time instead of 3 or 4 we have 1/3 raised to 'x', so first get me the graph of \(\large (\frac 1 3)^x\), the same way as before.

OpenStudy (anonymous):